OpenStudy (anonymous):

need help-asap plz; x *x+x=-25 what is the value of x.

OpenStudy (anonymous):

x*x+x=-25 x^2+x=-25 add +25 for both sides x^2+x+25=-25+25=0 \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

b=1 a=1 c=25

OpenStudy (anonymous):

(-1+-sqrt(1-4*1*25))/2*1 -1/2+- sqrt(1-100)/2 -1/2+-sqrt(-99)/2